Exercise 22-16 (Algorithmic) (LO. 10) TyroneCo, an S corporation with a positive E & P balance,...

  1. Exercise 22-16 (Algorithmic) (LO. 10)

    TyroneCo, an S corporation with a positive E & P balance, reports gross receipts for the year totaling $453,000 (of which $226,500 is passive investment income). Expenditures directly connected to the production of the passive investment income total $90,600.

    Compute Tyrone’s passive investment income tax.
    If required, round your final answer to the nearest dollar.
    $
Answer #1

Excessive net passive income = passive investment income in excess of 25% of gross receipts for the year / passive investment income for the year x net passive investment income for the year: 453,000-(453,000×25%)/453,000 ×135,900(226,500-90,600)

=$101,925.

Passive investment income tax = $101,925 ×35%

= $35,674.

____×____

All the best
Answer #2

Problem 12-21 (LO. 10)

TyroneCo, an S corporation with a positive E & P balance, reports gross receipts for the year totaling $400,000 (of which $200,000 is passive investment income). Expenditures directly connected to the production of the passive investment income total $80,000.

Compute Tyrone's passive investment income tax.
If required, round your final answer to the nearest dollar.
$

1. 20 0000-80 000=120 000

2. 400 000-25%=100 000; 200 000-100 000= 100 000

3. 100 000 : 200 000 = X 120 000 = 60 000 X 21% = 12 600  ;D


answered by: anonymous